# Mamba Neural Operator: Who Wins? Transformers vs. State-Space Models for PDEs

### v2

@@ III-E Mamba for Neural Operators
$i_{t+1}(x):=\sigma\left(Wi_{t}(x)+K_{\phi}(a)i_{t}(x)\right),\quad\forall x\in D,$ (13)
@@
(Kernel integral operator $K$): Define the kernel integral operator mapping in 2 by
[…]
@@
As mentioned in the previous section, we can be discrete SSM into the form of (7). This representation can be rewrite as[33]: $\mathbf{h}_{b}=\mathbf{w}_{T}\odot\mathbf{h}_{a}+\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i}).$ We define $\mathbf{V}=[\mathbf{V}_{1};\ldots;\mathbf{V}T]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{V}_{i}=u_{a+i-1}\Delta_{a+i-1}\in\mathbb{R}^{1\times D_{v}}$, $\mathbf{K}=[\mathbf{K}_{1};\ldots;\mathbf{K}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{K}_{i}=\mathbf{B}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$ , and $\mathbf{Q}=[\mathbf{Q}_{1};\ldots;\mathbf{Q}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{Q}_{i}=\mathbf{C}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$. We further define $\mathbf{w}=[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\Delta_{a-1+j}}\in\mathbb{R}^{D_{k}\times D_{v}}$, and $\mathbf{H}=[\mathbf{h}_{a};\ldots;\mathbf{h}_{b}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{h}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$. Finally, we set $\mathbf{Y}=[\mathbf{y}_{a};\ldots;\mathbf{y}_{b}]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{y}_{i}\in\mathbb{R}^{D_{v}}$
This formulation indicates that Gated Linear Attention[34] is actually a specific variant of Mamba. We next present our main result is how neural operator layers share a comparable structural framework with time-varying SSMs, which, to the best of our knowledge, is established here for the first time.
@@
$K_{\phi}(a)i_{t}(x):=\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i})$ (17)
@@
[…]
IV Experiments and Discussion
@@
IV-A Dataset Description & Implementation Protocol
PDEs Selection. We utilise datasets from PDEBench[35], a publicly available benchmark for partial differential equations (PDEs). We focus on three PDEs representing both stationary and time-dependent problems: Darcy Flow, Shallow Water 2D (SW2D), and Diffusion Reaction 2D (DR2D). All simulations are performed on a uniform grid. Detailed information about the datasets is defined as follows:
@@ IV-A Dataset Description & Implementation Protocol
with $a(x,y)$ and $u(x,y)$ as previously defined. Similar as PDEBench[35] protocol, we used only $\beta=1.0$ and we divided the training and testing ratio into 9:1 which contains 9,000 samples for training and 1,000 samples for testing.
@@
Each sample in the dataset is discretized on a spatial grid of $128^{2}$ points and a temporal grid of 101 time steps. The first 10 time steps are used as input to the model, while the remaining 91 time steps serve as the target output. Following the protocol established by PDEBench[35], the dataset consists of 900 samples for training and 100 samples for testing.
@@
where $t^{\prime}=0.045\,\text{s}$ and $T=5.000\,\text{s}$, and the spatial domain is $\Omega=[-1,1]^{2}$. Here, $u=u(x,y,t)$ and $v=v(x,y,t)$ represent the activator and inhibitor, respectively. In this dataset, we follow the same discretization scheme similar to the Shallow Water equation, where each sample is downsampled to a spatial resolution of $128^{2}$ and a temporal resolution of 101 time steps (with 10 for input and rest of the 91 for target). Similar as the PDEBench protocol[35], the dataset includes 900 samples for training and 100 samples for testing.
@@
Implementation & Evaluation Protocol. As Transformers have become the go-to architecture for PDE modelling and serve as the primary counterpart to SSM models, we selected three state-of-the-art Transformers as our baselines: GNOT[29], Galerkin Transformer (G.T.)[9], and OFormer[10]. To achieve a fair comparison between Transformers and Mamba, we integrated the S6 block and Cross S6 block to replace self-attention and cross-attention in each model, creating modified versions of the original architectures. All three experimental methods initially adopt a linear attention mechanism as described in their original publications, while we evaluated two configurations for each of them: an implementation with standard softmax attention mechanism (w/S.A.) and a Mamba-enhanced implementation (our Mamba Neural Operator principle) (w/Mamba). All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory to ensure consistent and fair comparison conditions. Three metrics including Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2) were utilised for evaluation.
@@
TABLE I: Quantitative comparison on Darcy Flow ($\beta=1$) across three methods with linear attention (original version), softmax attention and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted.
@@
GNOT[29] Galerkin 0.0070 0.0485 0.0370
@@
w/Mamba (MNO) Mamba 0.0061 0.0382 0.0286
OFormer[10] Normalised 0.0054 0.0253 0.0242
@@
TABLE II: Quantitative comparisons on Shallow Water 2D (SW2D) and Diffusion Reaction 2D (DR2D) across three methods with linear attention (original version) and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted in green.
SW2D DR2D
@@
GNOT[29]
Galerkin
0.0026 0.0025 0.0027 0.0567 0.6953 0.7233
w/Mamba (MNO) Mamba 0.0023 0.0022 0.0024 0.0060 0.0811 0.0570
G.T.[9]
Galerkin
0.0037 0.0035 0.0038 0.0083 0.1259 0.0723
w/Mamba (MNO) Mamba 0.0013 0.0013 0.0014 0.0012 0.0183 0.0099
OFormer[10] Normalised 0.0020 0.0020 0.0021 0.0177 0.2681 0.1559
w/Mamba (MNO) Mamba 0.0021 0.0021 0.0022 0.0123 0.1712 0.1134
IV-B Chose Your Winner: Transformer vs. Mamba for PDEs
We begin by evaluating the performance of Transformers, their variants, and Mamba on the Darcy Flow dataset, as presented in Table I. The results demonstrate that incorporating Mamba consistently improves performance across all metrics and models. For GNOT, while the RMSE remains close, the nRMSE and RL2 values are reduced, indicating that Mamba effectively refines predictions. The G.T. sees the most significant enhancement, with the RMSE dropping by 40% when Mamba is used. This suggests that Mamba’s design addresses the shortcomings of traditional Galerkin-type attention in capturing complex PDE dynamics. For OFormer, Mamba not only retains the strong baseline performance but also achieves improvements across all metrics. The reduction in RL2 indicates that Mamba’s mechanism is better at mapping the solution space of PDEs with higher precision. Mamba also demonstrates an enhanced ability to capture the complex spatial correlations inherent to Darcy Flow more effectively.
@@
Figure 4: Visualised prediction on Shallow Water dataset using Galerkin Transformer (G.T.) across the original and Mamba version.
Figure 5: Visualised prediction on Diffusion Reaction dataset using Galerkin Transformer (G.T.) across the original and Mamba version.
On the SW2D dataset, Mamba consistently outperforms the original Transformer models across all metrics. GNOT with Mamba achieves a lower RMSE and RL2, demonstrating Mamba’s ability to capture complex flow dynamics. The G.T. shows the most significant improvement, with RMSE of 65% reduction—highlighting Mamba’s superior capability in accurately representing the system’s behaviour. For OFormer, Mamba maintains comparable values but increases the RL2. On the DR2D dataset, the Mamba-enhanced models exhibit even more substantial gains. The G.T. sees a dramatic reduction in RMSE and RL2, showing Mamba’s strength in handling the complex dynamics.
@@
IV-C Why the Winner Wins: Breaking Down Mamba’s Win

### v3

@@ 3.5 Mamba for Neural Operators
$i_{t+1}(x)=\sigma\left(Wi_{t}(x)+K_{\phi}(a)i_{t}(x)\right),\quad\forall x\in D,$ (17)
Definition 3.
$i_{t+1}(x)=\sigma\left(Wi_{t}(x)+K_{\phi}(a)i_{t}(x)\right),\quad\forall x\in D,$ (18)
@@
Figure 2: Comparison of receptive fields and computational complexity across architectures. CNNs capture only local context with linear complexity, while ViTs achieve global modeling but with quadratic complexity. Our proposed Mamba (2D) attains global receptive fields with linear complexity by leveraging a bi-scan mechanism, combining efficiency with expressive global context.
@@
(Kernel integral operator $K$): Define the kernel integral operator mapping in 3 by
$K_{\phi}(a)i_{t}(x)=\int_{D}\kappa_{\phi}(x,y,a(x),a(y))i_{t}(y)\,dy,\quad\forall x,$ (19)
@@
As mentioned in the previous section, we can be discrete SSM into the form of (10). This representation can be rewrite as[liu2024vmamba]: $\mathbf{h}_{b}=\mathbf{w}_{T}\odot\mathbf{h}_{a}+\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i}).$ We define $\mathbf{V}=[\mathbf{V}_{1};\ldots;\mathbf{V}_{T}]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{V}_{i}=u_{a+i-1}\Delta_{a+i-1}\in\mathbb{R}^{1\times D_{v}}$, $\mathbf{K}=[\mathbf{K}_{1};\ldots;\mathbf{K}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{K}_{i}=\mathbf{B}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$ , and $\mathbf{Q}=[\mathbf{Q}_{1};\ldots;\mathbf{Q}_{T}]\in\mathbb{R}^{T\times D_{k}}$, where $\mathbf{Q}_{i}=\mathbf{C}_{a+i-1}\in\mathbb{R}^{1\times D_{k}}$. We further define $\mathbf{w}=[\mathbf{w}_{1};\ldots;\mathbf{w}_{T}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{w}_{i}=\prod_{j=1}^{i}e^{\mathbf{A}\Delta_{a-1+j}}\in\mathbb{R}^{D_{k}\times D_{v}}$, and $\mathbf{H}=[\mathbf{h}_{a};\ldots;\mathbf{h}_{b}]\in\mathbb{R}^{T\times D_{k}\times D_{v}}$, where $\mathbf{h}_{i}\in\mathbb{R}^{D_{k}\times D_{v}}$. Finally, we set $\mathbf{Y}=[\mathbf{y}_{a};\ldots;\mathbf{y}_{b}]\in\mathbb{R}^{T\times D_{v}}$, where $\mathbf{y}_{i}\in\mathbb{R}^{D_{v}}$
This formulation indicates that Gated Linear Attention[yang2023gated] is actually a specific variant of Mamba. We next present our main result is how neural operator layers share a comparable structural framework with time-varying SSMs, which, to the best of our knowledge, is established here for the first time.
@@
$K_{\phi}(a)i_{t}(x)=\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i})$ (22)
@@
According to Definition 4, it suffices to demonstrate that: $\int_{D}\kappa_{\phi}(x,y,a(x),a(y))i_{t}(y)\,dy=\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i})$. We assume the kernel $\kappa_{\phi}$ can be decomposed into a finite sum of separable basis functions: $\kappa_{\phi}(x,y,a(x),a(y))=\sum_{i=1}^{T}\omega_{i}\varphi_{i}(x)\psi_{i}(y)$ such that $\omega_{i}$ is learnable weights for each basis function. and Basis functions capturing interactions between $x$ and $y$. Then we substitute it into the integral such that : $\int_{D}\sum_{i=1}^{T}\omega_{i}\varphi_{i}(x)\psi_{i}(y)v_{t}(y)\,dy=\sum_{j=1}^{T}\omega_{i}\int_{D}\varphi_{i}(x)\psi_{i}(y)v_{t}(y)\,dy$. We further discretise the domain $D$ into $T$ points $\left\{\mathbf{y}_{i}\right\}_{i=1}^{T}$ with corresponding weights $\Delta y$. The integral becomes $K_{\phi}(a)i_{t}(\mathbf{x})\approx\sum_{i=1}^{T}\omega_{i}\sum_{j=1}^{T}\varphi_{i}(x)\psi_{i}(y_{j})i_{t}(\mathbf{y}_{j})\Delta y$. We represent $\varphi_{i}(\mathbf{x})$ as vector $\mathbf{K}_{i}$ and the input $i_{t}(\mathbf{y}_{i})$ as vector $\mathbf{V}_{i}$: $\mathbf{K}_{i}=\left[\varphi_{i}(\mathbf{x}),\varphi_{i}(\mathbf{x}),\dots,\varphi_{i}(\mathbf{x})\right]^{\top}\in\mathbb{R}^{1\times D_{k}},\quad\mathbf{V}_{i}=\left[\psi_{i}(y_{1})i_{t}(y_{1})\Delta y,\dots,\psi_{i}(y_{T})i_{t}(y_{T})\Delta y\right]^{\top}\in\mathbb{R}^{1\times D_{k}}$. If we further factorise $\omega_{i}$ as $\frac{w_{T}}{w_{i}}$, where $w_{T}$ is a hyperparameter and $w_{i}$ represents a set of parameters to be learned, we obtain the update: $K_{\phi}(a)i_{t}(x):=\sum_{i=1}^{T}\frac{\mathbf{w}_{T}}{\mathbf{w}_{i}}\odot(\mathbf{K}_{i}^{\top}\mathbf{V}_{i})$. Consequently, the neural operator layer shares a comparable structural framework with time-varying SSMs, demonstrating that the hidden space update in these models aligns with the iterative process in neural operator layers.
FNO and DeepONet are known for their universal approximation properties and discretisation invariance. These properties enable them to generalise well across various PDE discretisations, but they do not specifically account for the connection between the model and the discretisation of the PDE being solved. In contrast, MNO is built on state-space models (SSMs) [doyle2013feedback], naturally formulated in continuous time, which provide an alignment to discretisation of the PDE. This allows MNO to leverage the structure of the underlying system in a way that FNO and DeepONet do not, ensuring better alignment with the problem-specific discretisation.
The state-space formulation enables explicit control over discretisation-induced errors [basar2001new], providing theoretical guarantees on convergence and stability that are not inherently available in universal approximator-based methods like FNO and DeepONet. Furthermore, while FNO and DeepONet can learn the solution operator through deep neural networks or Fourier-based methods, these approaches do not explicitly model the temporal dynamics or interactions within the underlying PDE system in the same way MNO does. By framing the problem in terms of a state-space system, MNO provides a more robust and interpretable framework, especially when scaling to higher resolutions or dealing with complex boundary conditions, which are common in real-world PDE applications.
4 Experiments and Discussion
@@
4.1 Dataset Description & Implementation Protocol
PDEs Selection. We utilise datasets from PDEBench[takamoto2022pdebench], a publicly available benchmark for partial differential equations (PDEs). We focus on three PDEs representing both stationary and time-dependent problems: Darcy Flow, Shallow Water 2D (SW2D), Diffusion Reaction 2D (DR2D), and Compressible Navier-Stokes equation 2D (CFD2D). All simulations are performed on a uniform grid. We also included ND2D-c which has a irrengular mesh and is multi-scale. Detailed information about the datasets is defined as follows:
@@ 4.1 Dataset Description & Implementation Protocol
with $a(x,y)$ and $u(x,y)$ as previously defined. Similar as PDEBench[takamoto2022pdebench] protocol, we used only $\beta=1.0$ and we divided the training and testing ratio into 9:1 which contains 9,000 samples for training and 1,000 samples for testing.
@@
Each sample in the dataset is discretised on a spatial grid of $128^{2}$ points and a temporal grid of 101 time steps. The first 10 time steps are used as input to the model, while the remaining 91 time steps serve as the target output. Following the protocol established by PDEBench[takamoto2022pdebench], the dataset consists of 900 samples for training and 100 samples for testing.
@@
where $t^{\prime}=0.045\,\text{s}$ and $T=5.000\,\text{s}$, and the spatial domain is $\Omega=[-1,1]^{2}$. Here, $u=u(x,y,t)$ and $v=v(x,y,t)$ represent the activator and inhibitor, respectively. In this dataset, we follow the same discretization scheme similar to the Shallow Water equation, where each sample is downsampled to a spatial resolution of $128^{2}$ and a temporal resolution of 101 time steps (with 10 for input and rest of the 91 for target). Similar as the PDEBench protocol[takamoto2022pdebench], the dataset includes 900 samples for training and 100 samples for testing.
2D Compressible Navier-Stokes equation.
The compressible fluid dynamics equations govern fluid flows, expressed as:
$\partial_{t}\rho+\nabla\cdot(\rho\mathbf{v})=0,$ (31)
$\rho\left(\partial_{t}\mathbf{v}+\mathbf{v}\cdot\nabla\mathbf{v}\right)=-\nabla p+\eta\Delta\mathbf{v}+\left(\zeta+\frac{\eta}{3}\right)\nabla(\nabla\cdot\mathbf{v}),$ (32)
$\partial_{t}\left(\epsilon+\frac{\rho v^{2}}{2}\right)+\nabla\cdot\left[\left(p+\epsilon+\frac{\rho v^{2}}{2}\right)\mathbf{v}-\mathbf{v}\cdot\boldsymbol{\sigma}^{\prime}\right]=0,$ (33)
where $\rho$ is the mass density, $\mathbf{v}$ denotes the fluid velocity, $p$ is the thermodynamic pressure, and $\epsilon$ represents the internal energy density determined through an equation of state. The tensor $\boldsymbol{\sigma}^{\prime}$ is the viscous stress tensor, while $\eta$ and $\zeta$ correspond to the shear and bulk viscosity coefficients. This equation can capture complex phenomena, including the formation and propagation of shock waves, and finds applications in diverse real‐world scenarios such as aerodynamic flow around aircraft wings and the dynamics of interstellar gases. For more details of the dataset, we refer to [takamoto2022pdebench] Appendix D.
NS2D-c.
This problem involves a 2D steady-state Navier–Stokes system posed on a rectangular domain with four circular holes. Specifically, the domain is $\Omega=[0,8]^{2}\setminus\bigcup_{i=1}^{4}R_{i}$, where $R_{i}$ denotes the $i$-th circular region. The governing equations are
$(u\cdot\nabla)u$ $=\frac{1}{\mathrm{Re}}\nabla^{2}u-\nabla p,$ (14)
$\nabla\cdot u$ $=0,$ (15)
where $u$ is the velocity field, $p$ is the pressure, and $\mathrm{Re}$ is the Reynolds number. At the outlet, the pressure is fixed at zero, while at the inlet, the horizontal velocity component is prescribed as $u_{x}=\frac{y(8-y)}{16}$.
@@
Implementation & Evaluation Protocol. As Transformers have become the go-to architecture for PDE modelling and serve as the primary counterpart to SSM models, we selected three state-of-the-art Transformers as our baselines: GNOT[hao2023gnot], Galerkin Transformer (G.T.)[cao2021choose], and OFormer[li2022transformer]. In addition, we also included several non-transformer-based baselines: FNO [li2020fourier] , UNet [ronneberger2015u], DeepONet [lu2019deeponet] To achieve a fair comparison between Transformers and Mamba, we integrated the S6 block and Cross S6 block to replace self-attention and cross-attention in each model, creating modified versions of the original architectures. All three experimental methods initially adopt a linear attention mechanism as described in their original publications, while we evaluated two configurations for each of them: an implementation with standard softmax attention mechanism (w/S.A.) and a Mamba-enhanced implementation (our Mamba Neural Operator principle) (w/Mamba). All experiments were conducted on a single NVIDIA RTX 4090 GPU with 24GB of memory to ensure consistent and fair comparison conditions. Three metrics including Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2) were utilised for evaluation. For the testing time experiments, we take the average of 100 run on NVIDIA RTX 4090 GPU with 24GB of memory and Intel Xeon w7-3455.
@@
Table 1: Quantitative comparison on Darcy Flow ($\beta=1$) across three methods with linear attention (original version), softmax attention and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best-performing results highlighted.
@@
FNO (Li et al.2020) Non-Transformer 0.0119 0.0642 0.0363
UNet (Ronneberger et al.2015) Non-Transformer 0.0064 0.0325 0.0196
DeepONet (Lu et al.2019) Non-Transformer 0.0193 0.1023 0.1023
GNOT (Hao et al, 2023) Galerkin 0.0070 0.0485 0.0370
@@
w/Mamba (MNO) Mamba 0.0059 0.0374 0.0277
OFormer (Li et al, 2023) Normalised 0.0054 0.0253 0.0242
@@
Table 2: Quantitative comparisons on Shallow Water 2D (SW2D) and Diffusion Reaction 2D (DR2D) across three methods with linear attention (original version) and Mamba. The performance is measured in terms of Root Mean Squared Error (RMSE), Normalised RMSE (nRMSE), and Relative L2 Norm (RL2), with the best‐performing results highlighted in green.
SW2D DR2D CFD2D
@@
FNO Non-Transformer 0.0046 0.0044 0.0048 0.0081 0.1203 0.0715 - - -
UNet Non-Transformer 0.0858 0.0824 0.0914 0.0615 0.8449 0.5614 - - -
DeepONet Non-Transformer 0.0027 0.0026 0.0027 0.0667 0.8771 0.6040 - - -
GNOT Galerkin 0.0026 0.0025 0.0027 0.0567 0.6953 0.7233 3.9873 0.2447 0.3457
MNO Mamba 0.0023 0.0022 0.0024 0.0060 0.0811 0.0570 3.5681 0.2309 0.3113
G.T. Galerkin 0.0037 0.0035 0.0038 0.0083 0.1259 0.0723 4.4961 0.8806 0.6380
MNO Mamba 0.0013 0.0013 0.0014 0.0012 0.0183 0.0099 0.5014 0.4813 0.0736
OFormer Normalised 0.0020 0.0020 0.0021 0.0177 0.2681 0.1559 0.5181 0.5938 0.0973
MNO Mamba 0.0021 0.0021 0.0022 0.0123 0.1712 0.1134 0.5018 0.5897 0.0889
4.2 Choose Your Winner: Transformer vs. Mamba for PDEs
We begin by evaluating the performance of Transformers, their variants, and Mamba on the Darcy Flow dataset, as presented in Table 1. The results demonstrate that incorporating Mamba consistently improves performance across all metrics and models. Among the non-Transformer baselines (FNO, UNet, DeepONet), UNet is the strongest performer. Compared to this best baseline, our Mamba-based models still secure a further  15.6% reduction in RMSE and  25% reduction in nRMSE, highlighting Mamba’s advantage, even outperforming non-transformer methods. For GNOT, while the RMSE remains close, the nRMSE and RL2 values are reduced, indicating that Mamba effectively refines predictions. The G.T. sees the most significant enhancement, with the RMSE dropping by 40% when Mamba is used. This suggests that Mamba’s design addresses the shortcomings of traditional Galerkin-type attention in capturing complex PDE dynamics. For OFormer, Mamba not only retains the strong baseline performance but also achieves improvements across all metrics. The reduction in RL2 indicates that Mamba’s mechanism is better at mapping the solution space of PDEs with higher precision. Mamba also demonstrates an enhanced ability to capture the complex spatial correlations inherent to Darcy Flow more effectively.
@@
Figure 5: Visualized predictions on the Shallow Water dataset using the original Galerkin Transformer (G.T.) and our Mamba-enhanced version. The first row shows the ground truth, the second and third rows display the model predictions, and the final two rows present the corresponding absolute contour errors.
Figure 6: Visualized predictions on the Diffusion Reaction dataset using the original Galerkin Transformer (G.T.) and our Mamba-enhanced version. The first row shows the ground truth, the second and third rows display the model predictions, and the final two rows present the corresponding absolute contour errors.
On the SW2D dataset, Mamba consistently outperforms the original Transformer models across all metrics. Among the non-Transformer baselines (FNO, UNet, DeepONet), DeepONet is the strongest performer. Relative to this best baseline, our Mamba-based models still secure a further  51.9% reduction in RMSE and  50% reduction in nRMSE. GNOT with Mamba achieves a lower RMSE and RL2, demonstrating Mamba’s ability to capture complex flow dynamics. The G.T. shows the most significant improvement, with RMSE of 65% reduction—highlighting Mamba’s superior capability in accurately representing the system’s behaviour. For OFormer, Mamba maintains comparable values but increases the RL2. On the DR2D dataset, the Galerkin-type Transformer (G.T.) equipped with Mamba (G.T. MNO) substantially outperforms all non-Transformer-based baselines. Compared to the best non-Transformer model, FNO, the G.T. MNO achieves a 85.2% reduction in RMSE (from 0.0081 to 0.0012), alongside similar proportional improvements in nRMSE and RL2. For Galerkin-based models, leveraging Mamba into GNOT yields a notable reduction in error, achieving a  89.4% decrease in RMSE and similar proportional improvements in nRMSE and RL2 compared to the original GNOT. The Galerkin Transformer (G.T.) also benefit a improvement , with Mamba reducing RMSE from 0.0083 to 0.0012, and RL2 from 0.0723 to 0.0099, marking a more than 80% improvement in RL2. For normalised attention models such as OFormer, Mamba further enhances accuracy, lowering RMSE from 0.0177 to 0.0123 and RL2 from 0.1559 to 0.1134. These results demonstrate that Mamba’s state-space formulation effectively captures the nonlinear spatio-temporal dependencies in DR2D, leading to both higher accuracy and improved stability over all tested architectures.
On the most challenging benchmark, CFD2D, trained at an exceptionally high spatial resolution of 512 $\times$ 512, our MNO consistently delivers the lowest errors and establishes a new state of the art. Comparing Transformer-based baselines—GNOT, a Galerkin Transformer (G.T.), and the OFormer—struggle to accurately capture the highly non-linear, vortical flow fields at this resolution. Re-implementing each of these architectures with the Mamba and bi-directional scan reduces their RMSE across the board: from 3.99 to 3.57 for GNOT (-10%), from 4.50 to 0.50 for the Galarkin Transformer (-89%), and from 0.518 to 0.502 for OFormer (-3%). Similar reductions appear in nRMSE and RL2, with the most dramatic improvement being an 88% reduction in RL2 for the G.T. baseline. The case of GNOT + Mamba, where the model achieves the lowest nRMSE but the highest RL2, indeed suggests that while the model performs well on average, it may struggle with capturing certain large-scale features. This discrepancy can arise from the fact that nRMSE tends to focus on relative error, which can mask larger deviations in regions with complex dynamics, while RL2 captures absolute errors across the domain, highlighting discrepancies in larger-scale structures. Importantly, at this high resolution, our MNO also surpasses all Transformer-based models, demonstrating its ability to scale effectively and retain superior accuracy even when solving computationally intensive PDEs on small grids. This underscores MNO’s capability to tackle highly challenging PDE tasks at high resolution.
@@
4.3 Why the Winner Wins: Breaking Down Mamba’s Win